OpenStudy (anonymous):

check my math pls lim as x ->infinite (x^2-x)/sqrt(x^4+x)

OpenStudy (anonymous):

i have this so far \[(x^2/x^2-x/x^2)/\sqrt(x^4/x^4+x/x^4)\]

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x^2-x}{\sqrt{x^4+x}}\] highest power is inside the sqrt( ) divide both to and bottom by (sqrt(x^4)) which you just did

OpenStudy (anonymous):

\[then (1-1/x)/\sqrt(1+1x^3)\]

OpenStudy (anonymous):

\[then (1-0)/\sqrt(1+0)\]

OpenStudy (anonymous):

and finally got 1

myininaya (myininaya):

that is exactly right \[\lim_{x \rightarrow \infty}\frac{\frac{x^2}{x^2}-\frac{x}{x^2}}{\sqrt{\frac{x^4}{x^4}+\frac{x}{x^4}}} =\lim_{x \rightarrow \infty} \frac{1-\frac{1}{x}}{\sqrt{1+\frac{1}{x^3}}}=\frac{1-0}{\sqrt{1+0}}=\frac{1}{\sqrt{1}}=\frac{1}{1}=1\]

OpenStudy (anonymous):

yep :D

OpenStudy (anonymous):

dont go yey i got a question

OpenStudy (anonymous):

let i have the same equation but instead of x^4 i have x^9 as denominator

OpenStudy (anonymous):

i would divide the top by x^3?

myininaya (myininaya):

so if we have \[\lim_{x \rightarrow \infty}\frac{x^2-x}{\sqrt{x^9+x}}\] divide top and bottom by sqrt(x^9)

myininaya (myininaya):

by sqrt(x^9) doesn't equal x^3

OpenStudy (anonymous):

what i meant was to divide the top by x^3

OpenStudy (anonymous):

but the bottom by x^9

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{\frac{x^2}{x^\frac{9}{2}}-\frac{x}{ x^\frac{9}{2}}}{\sqrt{\frac{x^9}{x^9}+\frac{x}{x^9}}} =\lim_{x \rightarrow \infty}\frac{\frac{1}{x^\frac{5}{2}}-\frac{1}{x^\frac{7}{2}}}{ \sqrt{ 1-\frac{1}{x^\frac{7}{2}}}} \]

OpenStudy (anonymous):

:O i see that 2 from x^9/2 is from the sqrt right? so if it was 3sqrt it would x^9/3 so just x^3

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x^2-x}{\sqrt[3]{x^9-x}} =\lim_{x \rightarrow \infty}\frac{\frac{x^2}{x^3}-\frac{x}{x^3}}{\sqrt[3]{\frac{x^9}{x^9}-\frac{x}{x^9}}}\] yea

OpenStudy (anonymous):

cool thanks :D